OpenStudy (anonymous):

Solve. x – 6 = –18 (1 point) Do I Add 18 And 6 Or Minus?

OpenStudy (anonymous):

Add six and the anwser is -12

OpenStudy (anonymous):

Wait What?

OpenStudy (anonymous):

x - 6 = -18 +6 +6 x = -12

OpenStudy (anonymous):

Oh Alright .